Monthly income, in thousands of dollars ($K), at the Acme Law Firm for the six-month period from October to March is listed in the table below.

month

income, $K

Oct

69.3

Nov

68.6

Dec

69.4

Jan

71.7

Feb

72.3

Mar

74.8

-what is the simple exponential smoothing forecast with  α = 0.1 for monthly income for April?

-what is the regression forecast for monthly income for April?

-what is the numerical value of the strength of the linear association between monthly income and month number?

-how much of the variation in monthly income is accounted for by the regression?

-what is the approximate average rate of change of monthly income?

-What forecast method is the most accurate?

what is the best forecast value for monthly income for April?
Answer #1

1)

Simple exponential smoothing forecast with  α = 0.1 for monthly income for April is computed below

A E 1 Month 2 Oct 3 Nov 4 Dec 5 Jan 6 Feb 7 Mar 8 Apr B Income $K 69.3 68.6 69.4 71.7 72.3 74.8 С Exponential smoothing forec

EXCEL FORMULA:

А B с D 1 Month Income $K Exponential smoothing forecast Absolute Deviation 2 Oct 69.3 =B2 =ABS(B2-C2) 3 Nov 68.6 =+B2*0.1+C2

Simple exponential smoothing forecast with  α = 0.1 for monthly income for April = 70.28

-----------------------------------------------

2)

Regression forecast

X 1 2 3 4 5 6 21 69.3 68.6 69.4 71.7 72.3 74.8 426.1 x*y 69.3 137.2 208.2 286.8 361.5 448.8 1511.8 x? 1 4 9 16 25 36 91 y 480

Regression forecast for monthly income for April = 66.9266+1.1686*7 = 75.11

-----------------------------------------------

3)

Numerical value of the strength of the linear association between monthly income and month number is indicated by correlation coefficient (r)

Coefficient of correlation, P = *Σ- (Σκ) Σν) (Σ)-(Σκ) (Σ)- (Σ») f= 0.9267

Value of r is 0.9267, which is positive and close to 1

Which means linear association between monthly income and month number is direct and  strong

-----------------------------------------------

4)

How much of the variation in monthly income is accounted for by the regression

This is indicated by coefficient of determination, r2

Coefficient of determination, r2 is the measure of percentage of variability of relationship between the two variables. r2 =

85.87 % variation in monthly income is accounted for by the regression

-----------------------------------------------